OpenStudy (anonymous):

One solution to the problem below is 3. What is the other solution? b^2-9=0

OpenStudy (phi):

do you know how to factor a "difference of squares" ?

OpenStudy (anonymous):

well square roots mean even a negative number will be positive -3^2 =9-9=0

OpenStudy (phi):

when you see (a^2 - b^2) try to remember you can factor that into (a+b)(a-b)

OpenStudy (phi):

the other way to answer is to add +9 to both sides b^2 -9+9 = 0+9 b^2 = 9 now take the square root of both sides \[ \sqrt{b^2} = \sqrt{9} \\ b= \pm 3 \]

OpenStudy (anonymous):

sissy i dont know..you know we both are dumb..thats why we both in summer school honey <3

OpenStudy (anonymous):

its -3

OpenStudy (anonymous):

there ya go sis ^

OpenStudy (anonymous):

thanks you get aaaaaaa Medal!!! ooooo